# Enhancing the Generalization Capability of Skin Lesion Classification Models with Active Domain Adaptation Methods

### v1

@@ 4.3 Model Training
When retraining the model with DINO SSL on the in-domain data, we use all the skin lesion datasets from all the domains. We use the default data augmentations, with random cropping, 2 bigger view sizes (0.4-1.0 scale) and 4 smaller view sizes (0.05-0.4 scale), rotation, jitter, blur and black and white transformations. Only the bigger view sizes are passed to the teacher model, while all the view sizes are passed to the student model. This is to encourage learning “local-to-global” correspondence. We have tried different data augmentations, and found that the default setting gives the best result. The training process we find working best is that first, we freeze the backbone model and only train the prediction head layers with a learning rate of 1e-3. We train for 30 epochs. Then we train all the layers with a learning rate of 5e-4. We train for 25 epochs at which point the loss plateaus.
@@
We then finetune the model on the skin lesion source domain dataset, H. We freeze the backbone and only train the classifier head which is a single linear layer, because we want to compare the quality of the backbone features.
Model training is done on a RTX 4090 GPU. The batch size is 128. The image size is 224 x 224.
4.4 Ablation Study
We do ablation study to find the benefits of each step in the workflow, specifically 1. SSL training. 2. Applying the UDA DANN method. 3. Applying ADA (random with finetuning vs. AADA).
5 Results and discussions
Figure 2: Performance comparison of SL and SSL training.
First we compare the generalization capability of a SL pretrained model on imagenet vs. a SSL pretrained model on imagenet. We compare their classification performance on target domains with the AUPRC metric. In Figure 2, we see that the DINO pretrained model gives better results on 8 of the 11 domains, except BLH, BA, and HA which have slightly worse performance. When comparing the DINO pretrained model on imagenet vs. DINO retraining on skin lesion datasets, we see that DINO retraining gives better results on 7 of the 11 domains and on other domains the results are similar. So DINO retraining on in-domain data is helpful.
Figure 3: Effect of applying UDA on different target domains.
In Figure 3, for most of the datasets, with and without UDA give a similar result. But with UDA, it mainly improves for the HLP dataset. When compared with the results reported in [9], our results are better than their UDA results on 6 of the 10 target domains, except on HA, BA, BLH, and MLH having slightly worse results. This shows that SSL pretraining is an effective UDA method, so applying UDA method doesn’t help much. Since it doesn’t decrease the performance and may increase for few datasets, we still apply the additional UDA method.
Figure 4: Effect of applying ADA for 10 iterations.
Here we apply 10 iterations of ADA with each iteration adding 10 samples for annotation. The baseline ADA method is using random selection and finetuning. We choose number 10, since it’s a small and practical number for dermatologists to work with. In Figure 4, we see generally the performance increase with the increasing amount of labels, with the exception of HA and HLP datasets. We see by 3 iterations of ADA with total 30 labels, the performance plateaus. Using AADA and random sampling have a similar performance, but AADA’s performance is more stable as iteration increases.
6 Conclusion
In this paper, we propose a workflow of improving skin lesion classification models’ generalization capability by combining SSL, UDA and ADA methods. Our results show the benefits of each step of the workflow. This workflow helps models to perform well in different domains which mimic the actual clinical use cases, where domain shifts commonly exist.
For future studies, different model architecture can be considered, e.g. vision transformers, since they are shown to give good performance on natural image datasets.

### v2

@@ IV-C Model Training
We obtained the pre-trained DINO Resnet50 model from Facebook Research’s GitHub repository  [^1] https://github.com/facebookresearch/dino?tab=readme-ov-file. When retraining the model with DINO SSL on the skin lesion data, we utilized all datasets listed in Table I. We employed the default DINO data augmentations, with random cropping, two larger view sizes (0.4-1.0 scale) and four smaller view sizes (0.05-0.4 scale), rotation, jitter, blur and black and white transformations. Only the larger view sizes are passed to the teacher model, whereas all the view sizes are passed to the student model. This encourages the model to learn “local-to-global” correspondence. We explored various data augmentation strategies, and found that the default setting yielded the best results. Our optimized training process involves initially freezing the backbone model and training only the prediction head layers with a starting learning rate (LR) of 1e-3. We trained for 30 epochs. Subsequently, we trained all the layers using adaptive learning rates, with initial LR values of 1e-4 for the backbone layers and 5e-4 for the projector layers. We employed the one-cycle LR scheduler. We trained for 25 epochs at which point the loss plateaued.
@@
We then fine-tuned the model on the skin lesion source domain dataset, H. We froze the backbone and trained only the classifier head (a single linear layer), to isolate the impact of the backbone feature quality.
Model training was performed on a RTX 4090 GPU. The batch size was 128, and the image size was 224 x 224 pixels.
V Results and discussions
We compared the skine lesion classification performance of different methods on target domains using the AUPRC metric. We present the comparison results alongside the baseline (no domain adaptation) and the BSP UDA method reported in [10]. Table II presents the results, with the mean and standard deviation reported across 5 random seeds.
First, compared with the results reported in [10] for the baseline and BSP methods, our approach achieved superior performance on 9 of the 10 target domains, with slightly worse results observed only for BLH dataset. This demonstrates that SSL retraining on in-domain data is an effective UDA method.
Next, we compared the performance of the SL pre-trained model with the SSL Dino pre-trained model on ImageNet. The SSL Dino pre-trained model could be further retrained on the skin lesion datasets. We observed that the SL pre-trained model performed worse than the SSL pre-trained model on 9 of the 10 datasets. Therefore, SSL pre-training is superior to SL pre-training for domain adaptation. Furthermore, the SSL Dino retrained model exhibited improved performance on 6 of the 10 datasets compared to the model without retraining. Thus, retraining on in-domain data is advantageous for domain adaptation.
Figure 4: Comparing ADA methods with AUPRC value.
Figure 5: Comparing ADA methods.
Here, we present the results of applying ADA methods with 10 annotated samples from the target domain. Figure 4 shows the AUPRC achieved by the ADA methods compared to the baseline (without ADA), which corresponds to the DINO retrained model from Table II.
We compared five different ADA methods. The ADA methods comprise two components: active learning and domain adaptation. The active learning methods we evaluated included CLUE, AADA, BADGE and uniform sampling. The domain adaptation methods we evaluated included MME [58], DANN, and fine-tuning. The simplest ADA method involves uniform (random) sampling and fine-tuning.
Significant performance variability exists across the different target domains, primarily due to the varying AUPRC baselines, which are directly related to the melanoma ratios. We observed performance variability for the same ADA methods across the target domains. This variability primarily stems from the domain shift relative to the source domain.
To better visualize the difference of the five ADA methods, Figure 5 presents the AUPRC delta between the ADA methods and the baseline. The AADA-DANN method exhibited the best performance. It performed better than or similar to the baseline on 9 of the 10 datasets, with the exception of the HLP dataset. One potential explanation is the small number of melanoma samples (only 15) in the HLP dataset, leading to greater variance in performance.
VI Conclusion
In this paper, we proposed a workflow to improve the generalization of skin lesion classification models by combining SSL and ADA methods. Our results demonstrated the benefits of each method. SSL proved to be an effective UDA method, and ADA provided further performance improvements. We compared five ADA methods and showed that AADA-DANN yielded the best performance. We evaluated the performance on 10 different datasets exhibiting varying degrees of domain shift. This mimics real-world clinical scenarios, where domain shifts are common. Thus, our approach can facilitate wider clinical adoption of the skin lesion classification models. Future studies could explore different backbone model architectures, such as vision transformers, given their demonstrated strong performance on natural image datasets.